### v1

# Sound Source Localization for Spatial Mapping of Surgical Actions in Dynamic Scenes

Abstract
Purpose: Surgical scene understanding is key to advancing computer-aided and intelligent surgical systems. Current approaches predominantly rely on visual data or end-to-end learning, which limits fine-grained contextual modeling. This work aims to enhance surgical scene representations by integrating 3D acoustic information, enabling temporally and spatially aware multimodal understanding of surgical environments.
Methods: We propose a novel framework for generating 4D audio-visual representations of surgical scenes by projecting acoustic localization information from a phased microphone array onto dynamic point clouds from a RGB-D camera. A transformer-based acoustic event detection module identifies relevant temporal segments containing tool–tissue interactions which are spatially localized in the audio-visual scene representation. The system was experimentally evaluated in a realistic operating room setup during simulated surgical procedures performed by experts.
Results: The proposed method successfully localizes surgical acoustic events in 3D space and associates them with visual scene elements. Experimental evaluation demonstrates accurate spatial sound localization and robust fusion of multimodal data, providing a comprehensive, dynamic representation of surgical activity.
Conclusion: This work introduces the first approach for spatial sound localization in dynamic surgical scenes, marking a significant advancement towards multimodal surgical scene representations. By integrating acoustic and visual data, the proposed framework enables richer contextual understanding and provides a foundation for future intelligent and autonomous surgical systems.
keywords:
Surgical Scene Understanding, Sound Source Localization, Multimodal Data, Acoustic Sensing
1 Introduction
Surgical procedures are inherently dynamic, involving complex interactions among surgical staff, patients, instruments, and other medical devices. Understanding these interactions — referred to as surgical scene understanding — will be a cornerstone of next-generation computer-aided surgery and a crucial enabler for intelligent and autonomous surgical systems capable of assisting or operating independently.
In surgical scene understanding research, a body of previous work proposed to explicitly model the relationships among surgeons, patients, and instruments using semantic and spatial information, e.g., as action triplets [1] or scene graphs [2], and use these intermediate representations for downstream tasks such as surgical phase or role prediction. In contrast to explicit modeling approaches, end-to-end training has recently demonstrated strong performance in predicting high-level surgical concepts (e.g., surgical phases) from low-level observations (e.g., individual video frames) [3, 4]. However, end-to-end learning-based approaches can be prone to learning non-causal relationships and dataset-specific biases [5]. Constructing a fine-grain digital representation of the surgical scene, often called a surgical digital twin, can be a promising framework to address this limitation. This resulting representation can then be used for geometric scene understanding in a subsequent stage, effectively decoupling low-level data processing from high-level semantic inference [5, 6]. While current approaches for the generation of digital surgery representations predominantly rely on visual data from optical camera systems, the integration of multimodal information has been shown to be beneficial for surgical scene understanding in end-to-end learning frameworks [3, 4] and is expected to play a pivotal role in advancing next-generation data-driven computer-aided surgery systems [7]. In this context, the integration of multimodal data in digital surgical scene representations has great potential to achieve a more fine-grain modeling of surgical procedures.
Recently, acoustic events generated during tool–tissue interactions—such as sawing, chiseling, or drilling—have been shown to provide valuable insights into surgical tasks [8, 9, 10, 11]. Furthermore, in our own previous work, we proposed a proof-of-concept for the localization of acoustic events in surgery based on a microphone array and acoustic beamforming in 2D image space [12]. The integration of acoustic information into digital surgical scene representations, however, requires spatial and temporal awareness to enable effective surgical scene understanding.
In this paper, we propose a novel paradigm to enable sound localization in dynamic surgical scenes, thereby enhancing digital surgical scene representations to achieve improved contextual understanding for intelligent surgical systems. The key contributions of this work can be summarized as follows:
  • We propose a novel concept to generate 4D audio-visual representations of surgical scenes by fusing dynamic point clouds obtained using a RGB-D camera with acoustic localization information obtained from an phased microphone array, also known as acoustic camera.
  • To identify relevant frames for event localization, we propose an acoustic event detection stage that is based on a transformer architecture and enables the detection of surgical acoustic events in a continuous surgical sequence to trigger the localization of detected acoustic events.
  • We thoroughly evaluate the proposed method with experimental data that has been recorded from a simulated surgical procedure executed by surgical experts in a realistic surgical environment.
To the best of the author’s knowledge, no comparable approach currently exists within the state of the art. The code and associated data will be made publicly available [^1] The code is available under https://github.com/matthiasseibold/ssl_v2. The associated data will be made available upon acceptance..
2 Related Work
To establish a systematic framework for describing and analyzing surgical procedures, previous research has introduced methods at different levels of abstraction. Examples include hierarchical context models of surgical tasks and sub-tasks within Surgical Process Models [13], as well as automated extraction of surgical phases from intraoperative video recordings [14]. Automated analysis and description of surgical workflows can support workflow optimization, facilitate automated documentation for reporting and education, detect complications or deviations from preoperative plans, and enable real-time alerts in the event of adverse occurrences [15].
More recently, action triplets have been proposed as an extension of high-level surgical phase models, explicitly capturing the interactions inherent to surgical procedures [1]. Furthermore, graph-based representations have been introduced that model surgical scenes as structured graphs, where nodes correspond to relevant actors and objects in the operating room and edges represent spatial and functional relationships between them [2, 16, 17]. State-of-the-art approaches in this area primarily rely on visual data acquired from RGB or RGB-D cameras to derive both semantic and spatial information for modeling the surgical context. While end-to-end learning-based approaches achieve state-of-the-art performance for surgical scene graph prediction and leverage multimodal information [3, 4], they have limitations such as lack of explainability and the potential to learn non-causal relationships [5].
Looking ahead, multimodal data will be crucial for achieving a more comprehensive and fine-grained understanding of surgical processes, addressing the limitations of unimodal and particularly visual systems [7]. Among these modalities, acoustic sensing has emerged as a promising approach due to its radiation-free, low-cost, non-destructive, and easily integrable nature. Acoustic signals can capture time-critical events (e.g., drill breakthrough during surgical drilling [10]), provide complementary measurements to optical tracking or medical imaging, e.g., assessing press-fit during orthopedic implant insertion [9] or detect breaches during pedicle screw insertion [11], and detect characteristic acoustic events that indicate different surgical phases through ambient room microphones [8]. Despite these advances, current methods for the generation of fine-grain surgical scene representations still lack the capability to localize and correlate acoustic events within the 4D spatio-temporal context of surgery.
3 Methodology
Figure 1: Conceptual overview of the proposed system with an example obtained from surgical chiseling. In this example, hammer events are detected by the Acoustic Event Detection stage, as described in Section 3.3. Detected events trigger the Acoustic Event Localization stage, which projects the acoustic heatmap onto the dynamic 3D scene representation and localizes the event. We utilize point clouds from a co-calibrated RGB-D camera to represent the scene in our experiments, as described in Section 3.4. An optical tracking system is used for the evaluation of the system to compare the result with the ground truth pose of the surgical instrument.
To create a 4D audio-visual surgical scene representation, our system integrates acoustic information from a phased microphone array, also know as acoustic camera, with an existing spatio-temporal scene representation, for example obtained from a RGB-D camera. As illustrated in Figure 1, we propose an acoustic event detection module to detect acoustic events such as the use of surgical instruments, and an event localization module to spatially localize detected events in in the generated multimodal scene representation. The multimodal scene representation, event detection module, and event localization module are described in Sections 3.2, 3.3 and 3.4.
3.1 Experimental Setup and Data Collection
Figure 2: Overview of the experimental setup. We use our camera setup including the Ring48 microphone array, the ZED 2i RGB-D camera, and the FusionTrack 500 tracking system (shown in the left frame) to capture sequences of chiseling, drilling, and sawing (shown in the right frame).
We collected sequences of typical actions that are associated with sound emissions from Total Knee Arthoplasty using the experimental setup depicted in Figure 2. We simulate surgical sawing, drilling, and chiseling on a plastic bone model including simulated soft tissue, conducted by two trained surgical experts with real surgical tools. For drilling and sawing, we used a Arthrex DrillSaw Max 600 (Arthrex Inc., Naples, USA) surgical-grade power tool. We recorded sequences of length 20s20\>s with a frame rate of 25 fps and a sample rate of 192kHz192\>kHz from the acoustic camera for the surgical actions, resulting in $N_{chiseling}=6$, $N_{sawing}=9$, and $N_{drilling}=5$ sequences with a total length of 400s400\>s.
For the presented experiments, we use the gfai tech Ring48 (gfai tech GmbH, Berlin, Germany). The system employs a phased microphone array combined with acoustic beamforming to generate two-dimensional heatmaps of acoustic activity. An integrated optical sensor allows the acoustic activity map to be superimposed onto an optical image. To obtain a dynamic visual 3D representation of the surgical scene, we utilize a state-of-the-art RGB-D camera, namely the ZED 2i (StereoLabs, San Francisco, USA), which captures both color and depth information from the surgical scene.
To evaluate event localization module, we additionally integrate a FusionTrack 500 optical tracking system (Atracsys, Puidoux, Switzerland), which tracks instrument poses with sub-millimeter accuracy via attached infrared-reflective sphere markers. The attached marker arrays are calibrated to the system using 3D scans of the tools obtained with a high-fidelity 3D scanner, the Artec Space Spider (Artec3D, Senningerberg, Luxembourg). Based on the tracked instrument poses, we compute the ground-truth 3d bounding boxes for event locations. The ground truth event timestamps for the event detection were labeled manually based on visual and acoustic inspection of the synchronized audio and video recordings.
System Calibration and Synchronization $\mid$ All devices are calibrated and synchronized prior to the experiments. We calibrate the intrinsic parameters of the ZED 2i and Ring48 optical cameras using OpenCV and their extrinsic parameters $T_{ac}^{ftk}$ and $T_{zed}^{ftk}$ relative to the tracking system following the approach of [18], as shown in Figure 2. To ensure accurate synchronization between all devices throughout the experiments, we utilize a RocSync synchronization device [19], a hybrid clock comprising a linear and binary counter in both the RGB and IR spectrum, which enables a sub-frame accurate synchronization based on timestamps visible in all image frames. The factory-calibration of the phased microphone array is retained. At $1920\times 1080\>px$ resolution, final reprojection errors after calibration and synchronization are $2.30\pm 2.12\>px$ (Ring48) and $1.52\pm 0.84\>px$ (ZED 2i).
3.2 Multimodal Dynamic Surgical Scene Representation
We generate a multimodal, spatio-temporal scene representation based on the fusion of the acoustic information with the dynamic 3D point cloud. We compute 2D acoustic heatmaps with a fixed size of $100\times 100\>px$ via time-domain beamforming, using the professional sound analysis software NoiseImage 4.16.0 (gfai tech GmbH, Berlin, Germany). The heatmaps are computed at a predefined distance from the phased microphone array using a fixed 2D scanning grid. Band-filtering is applied between $1$ and 5kHz5\>kHz for sawing and between $1$ and 10kHz10\>kHz for drilling to obtain clear heatmaps without artifacts from ambient sound. For chiseling, no filtering is applied. The acoustic heatmaps are fused with the 3D scene representation by a simple projection operation. We project the amplitude of the acoustic heatmap, normalized to the range $[0,1]$. Future extensions may improve this step by projecting more expressive acoustic information, for example in the form of neural implicit sound descriptors.
3.3 Acoustic Event Detection
Our proposed acoustic event detection stage operates directly on the raw audio captured by the acoustic camera. We resample the audio with a sample rate of SR=16kHzSR=16\>kHz samples per second, apply a sliding window with length $L_{W}=150\>ms$ and hop length $L_{hop}=20\>ms$ to the sequence, and compute mel spectrograms with $n_{mels}=128$ mel bins from each window. For event detection, we fine-tune a pretrained AudioSpectrogramTransformer (AST) [20] in a binary classification task to predict event presence using PyTorch 2.8.0. During training we apply audio data augmentations, i.e., gain variation, gaussian noise, gain transitioning, clipping distortion, time stretching, and pitch shifting. The outputs are converted to video timestamps by detecting transitions in model predictions, as described in detail in the evaluation in Section 4. For sawing and drilling, we export multiple acoustic image frames in regular intervals after a detected event to evaluate the localization stage, as described in Section 4.
3.4 Event Localization
The location of a detected sound source is approximated within a 3D bounding box based on the computed multimodal scene representation. We localize areas with high signal amplitude within the scene using weighted clustering. Specifically, we use DBSCAN [21] with a neighborhood radius r=30.0mmr=30.0\>$m$, a minimum cluster weight of $m=200.0$, and the acoustic signal amplitudes as per-point weights. A tight bounding box is computed around the cluster with the largest total weight. Last, we limit the bounding box extents to a predefined range, based on the instrument size and the acoustic characteristics of each action. For drilling and sawing, the sound originates from the power tool surface due to motor-induced vibrations. Therefore, we limit the extents of the predicted 3D bounding box based on the known physical dimensions of the respective instruments. In contrast, for chiseling, sound emission is concentrated at the contact edge resulting from collisions between the mallet/chisel and the chisel/bone anatomy. Thus, we define a fixed-size bounding box with an edge length of 5 cm centered on the chisel’s contact edge.
4 Evaluation
Acoustic Event Detection $\mid$ Table 1 illustrates the quantitative results for the proposed event detection method described in Section 3.3. All results have been obtained using k-fold cross validation with parameters $k_{cv,chiseling}=k_{cv,drilling}=3$ and $k_{cv,sawing}=2$. We detect an event as a change in sign from $y_{pred}=0$ to $y_{pred}=1$ in the model predictions, where $y_{pred}=0$ corresponds to no event detected and $y_{pred}=1$ corresponding to a detected event. We introduce a relaxed condition, that allows for an event to be detected within a 20ms20\>ms window of $j$ frames around the ground truth event to be counted as a true positive prediction. We set this parameter as $j_{chiseling}=1$, $j_{sawing}=3$, and $j_{drilling}=10$. The hard condition only accepts predictions in the correct frame according to the ground truth. Figure 3 shows qualitative results for full example recordings.
Table 1: Precision, Recall, and F1 under hard and relaxed conditions
Method Precision Recall F1
Hard Relaxed Hard Relaxed Hard Relaxed
Chiseling $0.943\pm 0.047$ $\mathbf{0.986}\pm\mathbf{0.020}$ $0.756\pm 0.153$ $\mathbf{0.914}\pm\mathbf{0.063}$ $0.833\pm 0.101$ $\mathbf{0.948}\pm\mathbf{0.040}$
Drilling $0.000\pm 0.000$ $\mathbf{0.727}\pm\mathbf{0.196}$ $0.000\pm 0.000$ $\mathbf{0.687}\pm\mathbf{0.193}$ $0.000\pm 0.000$ $\mathbf{0.654}\pm\mathbf{0.057}$
Sawing $0.200\pm 0.000$ $\mathbf{1.000}\pm\mathbf{0.000}$ $0.200\pm 0.000$ $\mathbf{1.000}\pm\mathbf{0.000}$ $0.200\pm 0.000$ $\mathbf{1.000}\pm\mathbf{0.000}$
(a) Chiseling
(b) Drilling
(c) Sawing
Figure 3: We illustrate qualitative results over one full example recording of length $L_{sequence}=20s$, obtained from the experiment described in Section 4, for each of the surgical actions analyzed within this work. The detections have been obtained using the relaxed condition. The red arrows on top are the ground truth events, the green arrows on the bottom are the predicted events.
Acoustic Event Localization $\mid$ To evaluate our event localization approach, we compute the 3D intersection over union of the predicted and ground-truth 3D bounding boxes. For drilling and sawing actions, the ground-truth bounding box is computed from the instrument mesh and tracked pose obtained from the marker-based tracking system. For chiseling sequences, we define a the ground-truth is defined as a bounding box with a side length of 5cm5\>cm centered on the chisel’s contact edge, matching the definition used during inference. The distribution of IoU scores is shown in Figure 4. In Table 2, we evaluate the localization recall rates under multiple of intersection over union thresholds ranging from $0.05$ to $0.40$. Further qualitative results are provided in the supplementary videos.
Table 2: Recall rates of the event localization. A successful event localization is defined by the 3D bounding box IoU above a defined threshold.
IoU Threshold Chiseling Drilling Sawing Average
0.05 0.79 0.95 0.91 0.88
0.10 0.76 0.86 0.73 0.77
0.20 0.68 0.39 0.31 0.46
0.40 0.37 0.00 0.01 0.13
Figure 4: Distribution of 3D bounding box IoU scores over a total of 20 recordings, grouped by surgical action.
5 Discussion
In this paper, we introduce a novel concept that enables the generation of spatio-temporal and audio-visual surgical scene representations by mapping acoustic features generated by surgical actions to a dynamic point cloud of the scene. In this context, the presented work is the first to map acoustic information into digital surgical scene descriptions and has great potential for the development of fine-grain and multimodal models of surgical activity. In future work, this concept will be especially valuable when the resulting digital dynamic representation is further enriched with additional data from other low-level data science and computer vision algorithms, i.e., human poses, hand tracking, tool tracking, object detection, etc. Finally, the resulting representation, that is often called a surgical digital twin in related literature, can then be used as an intermediate representation for further downstream applications and to predict high-level surgical concepts, such as surgical scene graphs, action triplets, phases, roles, or for other applications like automated reporting.
The event detection module showed promising results, especially for chiseling and sawing, where events could be detected within a maximum offset of one frame for chiseling (corresponding to at most 20ms20\>ms), and a maximum offset of three frames (corresponding to at most 60ms60\>ms) with high accuracy. Especially, for sharp burst events, such as chiseling, an accurate detection is crucial. For sawing, the recordings contained only one onset and then a continuing sawing action, as visible in the waveform plot in Figure 3 which results in only one event present per sequence. Therefore, a relaxed criterion for the detection of the onset is tolerable. The detection of sawing events presented a more challenging problem, where we tolerate a maximum offset of 10 frames for a detected event (corresponding to at most 200ms200\>ms) and obtain decreased performance compared to the other surgical actions that have been analyzed within this work. This decreased performance can mainly be attributed to the fact that the surgeons turned on the device in the beginning of the sequence and kept drilling into the plastic bone while the power tool was running, resulting in only very subtle sound differences between drilling in the air and drilling into the rather soft plastic bone. To mitigate this issue, beam steering or noise canceling could improve the results, which is, however, beyond the scope of the presented work.
The event localization module proved that the fusion of audio-visual information is an effective approach to detect regions of interest with minimal computational overhead. While this method can directly be employed to inform down-stream tasks, such as scene graph generation or instrument tracking, we also see potential as an initial filtering stage that allows more computationally intensive methods to focus solely on the regions of interest. Despite the simplicity of our approach, $77\%$ of the events were successfully localized with a 3D bounding box intersection over union of at least $0.1$. This accuracy is enough to generate semantic scene states, i.e., scene graphs for surgical context understanding. Due to the low spatial resolution of the acoustic camera the localization accuracy is limited, as indicated by the steep decrease in recall under stricter intersection over union thresholds. Note however that our predictions were compared to posed ground-truth bounding boxes, whereas our approach does not explicitly optimize the rotation of the estimated bounding boxes. We argue that for region of interest detection a coarse localization is an acceptable trade-off for low latency. For tasks with higher localization requirements, the accuracy could be improved by occlusion-aware fusion of acoustic features into the scene representation. Future extensions using beam-steering based on the 3D scene geometry and a projection of more informative acoustic features, for example in the form of neural implicit sound representations, could bring further enhancements.
The exploratory character of our experiments inherently comes with some limitations. While the presented approach is real-time capable in theory, our current implementation only supports offline processing due to limitations in the software and API of the acoustic camera. We employ a single-view point cloud as the dynamic 3D visual representation of the surgical scene, which is the basis for mapping the localized acoustic activity. While the ZED2i RGB-D camera produces high-quality point clouds, the preservation of fine details and completeness of the scene is limited. To obtain a higher-quality visual representation, future research should utilize multi-view reconstructions to reduce occlusions and investigate alternative scene representations such as dynamic gaussian splats.
Acoustic information has been shown to have great potential for the analysis of surgical activity in related work, however, acoustic signal analysis is inherently limited to actions that generate audible sounds. While this makes the approach highly effective in sound-rich fields like orthopedics, its application to other surgical sub-fields may be more challenging. However, the proof-of-concept presented in this work does not fully utilize the high sensitivity of the acoustic camera. Especially for hard-to-detect sounds, acoustic beam steering can be a highly promising solution to improve the sound quality for acoustics-based downstream tasks.
6 Conclusion
In summary, this work introduces a novel framework for surgical scene understanding that fuses visual and acoustic information into a unified 4D audio-visual digital representation. By combining RGB-D–based dynamic point clouds with sound localization information obtained from a phased microphone array, and by incorporating transformer-based acoustic event detection, our approach not only enhances contextual awareness in surgical environments but also lays the groundwork for future intelligent and autonomous systems capable of richer, data-driven surgical scene interpretation.
Acknowledgement
This work has been funded by the Swiss National Foundation under the funding scheme ”Spark” with grant number 228813. The experiments were supported by the OR-X - a Swiss national research infrastructure for translational surgery and associated funding by the University Hospital Balgrist. Imaging was performed with support of the Swiss Center for Musculoskeletal Imaging (SCMI). We furthermore thank Dr. Matthias Brechbühl from Norsonic Brechbühl AG and Carsten Hessenius, Andy Meyer, and Michael Markus Ackermann from GFaI e.V. and gfai tech GmbH for the support that made this work possible.