### v2

# Sound Source Localization for Spatial Mapping of Surgical Actions in Dynamic Scenes

Abstract
Purpose: Surgical scene understanding is key to advancing computer-aided and intelligent surgical systems. Current approaches predominantly rely on visual data or end-to-end learning, which limits fine-grained contextual modeling. This work aims to enhance surgical scene representations by integrating 3D acoustic information, enabling temporally and spatially aware multimodal understanding of surgical environments.
Methods: We propose a novel framework for generating 4D audio-visual representations of surgical scenes by projecting acoustic localization information from a phased microphone array onto dynamic point clouds from a RGB-D camera. A transformer-based acoustic event detection module identifies relevant temporal segments containing tool–tissue interactions which are spatially localized in the audio-visual scene representation. The system was experimentally evaluated in a realistic operating room setup during simulated surgical procedures performed by experts.
Results: The proposed method successfully localizes surgical acoustic events in 3D space and associates them with visual scene elements. Experimental evaluation demonstrates accurate spatial sound localization and robust fusion of multimodal data, providing a comprehensive, dynamic representation of surgical activity.
Conclusion: This work introduces the first approach for spatial sound localization in dynamic surgical scenes, marking a significant advancement towards multimodal surgical scene representations. By integrating acoustic and visual data, the proposed framework enables richer contextual understanding and provides a foundation for future intelligent and autonomous surgical systems.
keywords:
Surgical Scene Understanding, Sound Source Localization, Multimodal Data, Acoustic Sensing
1 Introduction
Surgical procedures are inherently dynamic, involving complex interactions among surgical staff, patients, instruments, and other medical devices. To enable the next generation of intelligent computer-aided surgery systems, capable of context-aware assistance or semi-autonomous operation, machines must possess a holistic understanding of this environment. This capability, known as surgical scene understanding, relies on creating digital representations that accurately mirror the physical reality of the operating theater.
Historically, the field has prioritized visual perception. A body of works proposed to explicitly model the surgical context based on RGB or RGB-D data, employing semantic structures such as action triplets [nwoye2020triplets] or scene graphs [oeszoy2022ssg] to capture relationships between surgeons, tools, and tissues. In contrast to explicit modeling approaches, end-to-end training has recently demonstrated strong performance in predicting high-level surgical concepts (e.g., surgical phases) from low-level observations (e.g., individual video frames) [Ozsoy_2025_CVPR, oezsoy2025egoexoregoexocentricoperatingroom]. Nevertheless, these visual-centric approaches face inherent limitations. Visual data is susceptible to occlusion, lighting variations, and often fails to capture the physical nature of tool-tissue interactions, such as the mechanical resistance of bone during sawing or the precise moment of a drill breakthrough.
Multimodal surgical perception has emerged as a promising approach to overcome the limitations of unimodal vision. Several works proposed digital twins of surgery, which integrate complementary data streams such as kinematics, inertial measurements and system events into a comprehensive digital representation of the surgical scene, effectively decoupling low-level sensing from high-level semantic reasoning [ding2024dt, Hein_2024_CVPR]. Within this multimodal landscape, audio represents a largely untapped but highly information-rich modality.
Surgical sounds provide high-frequency cues about physical interactions that are often invisible to cameras. Recently, acoustic events generated during tool–tissue interactions—such as sawing, chiseling, or drilling—have been shown to provide valuable insights into surgical tasks [seibold2022conditional, goossens2020acoustic, seibold2021realtime]. Furthermore, in our own previous work, we proposed a proof-of-concept for the localization of acoustic events in surgery based on a microphone array and acoustic beamforming in 2D image space [seibold2024ssl]. The integration of acoustic information into digital surgical scene representations introduces unique capabilities, such as high temporal resolution (sub-fps) and the ability to detect events that cannot be captured by vision alone (e.g., activity of power tools), but requires spatial and temporal integration with the visual representation to enable effective multimodal surgical scene understanding. To the best of the author’s knowledge, there is currently no comparable approach that integrates spatial sound into dynamic surgical scene representations.
In this paper, we propose a novel paradigm to enable sound localization in dynamic surgical scenes, thereby enhancing digital surgical scene representations to achieve improved contextual understanding for intelligent surgical systems. The key contributions of this work can be summarized as follows:
  • We propose a novel concept to generate 4D audio-visual representations of surgical scenes defined by a time-varying visual geometric 3D representation of a surgical scene that is enriched with localized acoustic events occurring at certain points of time or spanning over a defined period. Our approach achieves this by fusing dynamic point clouds obtained using a RGB-D camera with acoustic localization information obtained from a phased microphone array, also known as acoustic camera.
  • To identify relevant frames for event localization, we propose an acoustic event detection stage that is based on a transformer architecture and enables the detection of surgical acoustic events in a continuous surgical sequence to trigger the localization of detected acoustic events.
  • We thoroughly evaluate the proposed method with experimental data that has been recorded from a simulated surgical procedure executed by surgical experts in a realistic surgical environment.
The code and associated data are available under https://github.com/matthiasseibold/ssl_v2.
2 Related Work
The proposed approaches for surgical scene understanding can be categorized by the modalities used to construct the digital representation, namely purely visual approaches, acoustic sensing, and multimodal fusion. The majority of existing works relies on optical data such as RGB or RGB-D to derive both semantic and spatial information for modeling the surgical context. Early approaches focused on temporal and hierarchical modeling, using surgical process models [neumuth2017spm] to extract phases and steps from video data [padoy2019workflow]. More recently, spatial understanding has been enhanced through action triplets (instrument, action, target) [nwoye2020triplets] and surgical scene graphs, which model the operating room as a structured graph of actors and objects [oeszoy2022ssg, oeszoy2023labrador, hamoud2024stor]. While end-to-end visual models achieve strong performance on specific benchmarks [Ozsoy_2025_CVPR], they lack the ability to capture non-visual physical phenomena and are prone to learning dataset-specific visual biases [ding2024dt].
Audio has emerged as a powerful complementary modality due to its non-destructive, radiation-free, and low-cost nature, and its ability to capture high-frequency mechanical events at a sub-frame temporal resolution. Prior work has demonstrated the utility of audio for surgical phase recognition [seibold2022conditional], real-time error prevention via breakthrough detection [seibold2021realtime], or assessing the press-fit of orthopedic implants [goossens2020acoustic]. However, these methods typically treat audio as a global signal detached from the scene geometry. While our previous work introduced a proof-of-concept for 2D sound source localization in image space [seibold2024ssl], it did not account for the 3D depth and dynamic nature of the surgical environment.
The integration of multimodal data, such as vision, kinematics, language, and audio, is widely recognized as essential for next-generation computer-aided surgery [hein2017sds]. An early multimodal work by Weede et al. [weede2012workflow] fuses positional information, audio and video signals in a Bayes classifier for surgical phase recognition. In the wake of recently emerging vision-language models, a series of works fuse vision and language modalities for decision-making in surgical robotics [zargarzadeh2025decision] and workflow analysis [yuan2025learning]. Yet, current approaches generally lack the capability to localize acoustic sources. This limitation prevents the system from distinguishing between simultaneous events (e.g., two tools operating at once) or associating a specific sound with a specific instrument in a crowded scene. By projecting acoustic localization information from a phased microphone array onto dynamic point clouds, our proposed method bridges this gap, enabling the first spatially aware, multimodal representation of surgical activity.
3 Methodology
Figure 1: Conceptual overview of the proposed system with an example obtained from surgical chiseling. In this example, hammer events are detected by the Acoustic Event Detection stage, as described in Section 3.3. Detected events trigger the Acoustic Event Localization stage, which projects the acoustic heatmap onto the dynamic 3D scene representation and localizes the event. We utilize point clouds from a co-calibrated RGB-D camera to represent the scene in our experiments, as described in Section 3.4. An optical tracking system is used for the evaluation of the system to compare the result with the ground truth pose of the surgical instrument.
To create a 4D audio-visual surgical scene representation, our system integrates acoustic information from a phased microphone array, also know as acoustic camera, with an existing spatio-temporal scene representation, for example obtained from a RGB-D camera. As illustrated in Figure 1, we propose an acoustic event detection module to detect acoustic events such as the use of surgical instruments, and an event localization module to spatially localize detected events in in the generated multimodal scene representation. The multimodal scene representation, event detection module, and event localization module are described in Sections 3.2, 3.3 and 3.4.
3.1 Experimental Setup and Data Collection
Figure 2: Overview of the experimental setup. We use the Ring48 microphone array, the ZED 2i RGB-D camera, and the FusionTrack 500 tracking system (shown in the left frame) to capture sequences of chiseling, drilling, and sawing (shown in the middle frame). The right frame shows the viewpoint of the acoustic and RGB-D camera.
We collected sequences of typical actions that are associated with sound emissions from Total Knee Arthoplasty using the experimental setup depicted in Figure 2. We simulate surgical sawing, drilling, and chiseling on a plastic bone model including simulated soft tissue (SYNBONE AG, Zizers, Switzerland), conducted by two trained surgical experts with real surgical tools. For drilling and sawing, we used a Arthrex DrillSaw Max 600 (Arthrex Inc., Naples, USA) surgical-grade power tool. We recorded sequences of length 20s20\>s with a frame rate of 25 fps and a sample rate of 192kHz192\>kHz from the acoustic camera for the surgical actions, resulting in $N_{chiseling}=6$, $N_{sawing}=9$, and $N_{drilling}=5$ sequences with a total length of 400s400\>s.
For the presented experiments, we use the gfai tech Ring48 (gfai tech GmbH, Berlin, Germany). The system employs a phased microphone array with acoustic beamforming to generate two-dimensional heatmaps of acoustic activity. An integrated optical sensor allows the acoustic activity map to be superimposed onto an optical image. To obtain a dynamic visual 3D representation of the surgical scene, we utilize a state-of-the-art RGB-D camera, namely the ZED 2i (StereoLabs, San Francisco, USA), which captures both color and depth information from the surgical scene.
To evaluate event localization module, we additionally integrate a FusionTrack 500 optical tracking system (Atracsys, Puidoux, Switzerland), which tracks instrument poses with sub-millimeter accuracy via attached infrared-reflective sphere markers. The attached marker arrays are calibrated to the system using 3D scans of the tools obtained with a high-fidelity 3D scanner, the Artec Space Spider (Artec3D, Senningerberg, Luxembourg). Based on the tracked instrument poses, we compute the ground-truth 3d bounding boxes for event locations. The ground truth event timestamps for the event detection were labeled manually based on visual and acoustic inspection of the synchronized audio and video recordings.
System Calibration and Synchronization All devices are calibrated and synchronized prior to the experiments. We calibrate the intrinsic parameters of the ZED 2i and Ring48 optical cameras using OpenCV and their extrinsic parameters $T_{ac}^{ftk}$ and $T_{zed}^{ftk}$ relative to the tracking system following the approach of [hein_next-generation_2025], as shown in Figure 2. To ensure accurate synchronization between all devices throughout the experiments, we utilize a RocSync synchronization device [meyer2025rocsync], a hybrid clock comprising a linear and binary counter in both the RGB and IR spectrum, which enables a sub-frame accurate synchronization based on timestamps visible in all image frames. The factory-calibration of the phased microphone array is retained. At $1920\times 1080\>px$ resolution, final reprojection errors after calibration and synchronization are $2.30\pm 2.12\>px$ (Ring48) and $1.52\pm 0.84\>px$ (ZED 2i).
3.2 Multimodal Dynamic Surgical Scene Representation
We generate a multimodal, spatio-temporal scene representation based on the fusion of the acoustic information with the dynamic 3D point cloud. We compute 2D acoustic heatmaps with a fixed size of $100\times 100\>px$ via time-domain beamforming, using the professional sound analysis software NoiseImage 4.16.0 (gfai tech GmbH, Berlin, Germany). The heatmaps are computed at a predefined distance from the phased microphone array using a fixed 2D scanning grid. Band-filtering is applied between $1$ and 5kHz5\>kHz for sawing and between $1$ and 10kHz10\>kHz for drilling to obtain clear heatmaps without artifacts from ambient sound. For chiseling, no filtering is applied. The acoustic heatmaps are fused with the 3D scene representation by a simple projection operation. We project the amplitude of the acoustic heatmap, normalized to the range $[0,1]$.
3.3 Acoustic Event Detection
Our proposed acoustic event detection stage operates directly on the raw audio captured by the acoustic camera. We resample the audio with a sample rate of SR=16kHzSR=16\>kHz samples per second, apply a sliding window with length $L_{W}=150\>ms$ and hop length $L_{hop}=20\>ms$ to the sequence, and compute mel spectrograms with $n_{mels}=128$ mel bins from each window. For event detection, we fine-tune a pretrained AudioSpectrogramTransformer (AST) [gong21b_interspeech] in a multi-class classification task (classes = idle, chiseling, drilling, sawing) to predict event presence using PyTorch 2.8.0. During training we apply audio data augmentations, i.e., gain variation, gaussian noise, gain transitioning, clipping distortion, time stretching, and pitch shifting. The outputs are converted to video timestamps by detecting transitions in model predictions, as described in detail in the evaluation in Section 4. For sawing and drilling, we export multiple acoustic image frames in regular intervals after a detected event to evaluate the localization stage, as described in Section 4.
3.4 Event Localization
The location of a detected sound source is approximated within a 3D bounding box based on the computed multimodal scene representation. We localize areas with high signal amplitude within the scene using weighted clustering. Specifically, we use DBSCAN [ester1996density] with a neighborhood radius r=30.0mmr=30.0\>$m$, a minimum cluster weight of $\color[rgb]{0,0,0}w=400.0$}, and the acoustic signal amplitudes as per-point weights. A tight bounding box is computed around the cluster with the largest total weight. Last, we limit the side lengths of the predicted bounding box to a predefined range, based on the instrument size and the acoustic characteristics of each action. For drilling and sawing, the sound originates from the power tool surface due to motor-induced vibrations. Therefore, we limit the box side lengths to the minimum and maximum side length of the instrument’s ground-truth bounding box. In contrast, for chiseling, sound emission is concentrated at the contact edge resulting from collisions between the mallet/chisel and the chisel/bone anatomy. Thus, we define a fixed-size bounding box with an edge length of 5 cm centered on the chisel’s contact edge. Since this step is executed after the event detection, the approximate physical dimensions of the sound-emitting instruments can be assumed to be known.
4 Evaluation
Acoustic Event Detection Table 1 illustrates the quantitative results for the proposed event detection method described in Section 3.3. All results have been obtained using 3 -fold cross validation. We detect an event as a change in sign from $y_{pred}=0$ to $y_{pred}=1$ in the model predictions, where $y_{pred}=0$ corresponds to no event detected and $y_{pred}=1$ corresponding to a detected event. We introduce a relaxed condition, that allows for an event to be detected within a 20ms20\>ms window of $j$ frames around the ground truth event to be counted as a true positive prediction. We set this parameter as $j_{chiseling}=1$, $j_{sawing}=j_{drilling}=10$. The hard condition only accepts predictions in the correct frame according to the ground truth. Figure 3 shows qualitative results for full example recordings.
Table 1: Precision, Recall, and F1 under hard and relaxed conditions
Method Precision [%] Recall [%] F1 [%]
Hard Relaxed Hard Relaxed Hard Relaxed
Chiseling $0.915\pm 0.045$ $\mathbf{0.968}\pm\mathbf{0.022}$ $0.791\pm 0.134$ $\mathbf{0.953}\pm\mathbf{0.033}$ $0.845\pm 0.093$ $\mathbf{0.961}\pm\mathbf{0.026}$
Drilling $0.000\pm 0.000$ $\mathbf{0.753}\pm\mathbf{0.194}$ $0.000\pm 0.000$ $\mathbf{0.639}\pm\mathbf{0.160}$ $0.000\pm 0.000$ $\mathbf{0.651}\pm\mathbf{0.077}$
Sawing $0.611\pm 0.283$ $\mathbf{1.000}\pm\mathbf{0.000}$ $0.390\pm 0.079$ $\mathbf{0.889}\pm\mathbf{0.157}$ $0.444\pm 0.079$ $\mathbf{0.933}\pm\mathbf{0.094}$
(a) Chiseling
(b) Drilling
(c) Sawing
Figure 3: We illustrate qualitative results over one full example recording of length $L_{sequence}=20s$, obtained from the experiment described in Section 4, for each of the surgical actions analyzed within this work. The detections have been obtained using the relaxed condition. The red arrows on top are the ground truth events, the green arrows on the bottom are the predicted events.
Acoustic Event Localization To evaluate our event localization approach, we compute the 3D intersection over union of the predicted and ground-truth 3D bounding boxes. For drilling and sawing actions, the ground-truth bounding box is computed from the instrument mesh and tracked pose obtained from the marker-based tracking system. For chiseling sequences, we define a the ground-truth is defined as a bounding box with a side length of 5cm5\>cm centered on the chisel’s contact edge, matching the definition used during inference. The distribution of IoU scores is shown in Figure 4 where a IoU threshold of 0.1 results in recall rates of 0.78, 0.91, 0.84, and 0.84 for chiseling, drilling, sawing, and overall average, respectively. For further insights, please refer to the supplementary material, where we evaluate the localization recall rates under multiple of intersection over union thresholds ranging from $0.05$ to $0.40$. Further qualitative results are provided in the supplementary videos.
Figure 4: Distribution of 3D bounding box IoU scores and 3D bounding box center errors over a total of 20 recordings, grouped by surgical action.
4.1 Ablation Studies
To provide further insights into the result distribution, we perform a detailed analysis of the onset error distribution for the event detection stage of the proposed method and present the results in figure 5. Furthermore, we evaluate the robustness of the event detection stage under simulated realistic OR background noise in the supplementary material. In Figure 6, we analyze the sensitivity in the choice of DBSCAN parameters, as well as the density of the input point cloud. We vary one parameter at a time in the ranges of $r\in[7.5,60]\>mm$ and $w\in[50,800]$. In addition, we analyze the robustness to the input point cloud’s density by downsampling it by a factor of $s\in[\frac{1}{64},1]$.
Last, we compare our localization stage against a naive centroid detection baseline, which computes the 3D sound source location as the weighted average of the entire point cloud, using the sound amplitude as per-point weights. To obtain a region-of-interest bounding box, an axis-aligned bounding box of the instrument’s size is centered around the estimated location. This baseline achieves a 3D bounding box IoU of $0.14\pm 0.16$ compared to $0.23\pm 0.14$ of our localization stage. The 3D bounding box center error of this baseline is $144.10\pm 121.19\>mm$ compared to $101.39\pm 89.75\>mm$ obtained by the proposed localization stage.
Figure 5: The onset error distribution for the event detection component.
Figure 6: Sensitivity analysis of the DBSCAN radius and minimum weight parameters. We additionally evaluate the effect of the point cloud density on the localization accuracy by downsampling the input point cloud, while adjusting the minimum weight accordingly. All results are reported as the average IoU on the whole dataset.
5 Discussion
In this paper, we introduce a novel concept that enables the generation of spatio-temporal and audio-visual surgical scene representations by mapping acoustic features generated by surgical actions to a dynamic point cloud of the scene. In this context, the presented work is the first to map acoustic information into digital surgical scene descriptions and has great potential for the development of fine-grain and multimodal models of surgical activity. In future work, this concept will be especially valuable when the resulting digital dynamic representation is further enriched with additional data from other low-level data science and computer vision algorithms, i.e., human poses, hand tracking, tool tracking, object detection, etc. Finally, the resulting representation, that is often called a surgical digital twin in related literature, can then be used as an intermediate representation for further downstream applications and to predict high-level surgical concepts, such as surgical scene graphs, action triplets, phases, roles, or for other applications like automated reporting.
The event detection module showed promising results, especially for chiseling and sawing, where events could be detected within a maximum offset of one frame for chiseling (corresponding to at most 20ms20\>ms). Especially, for sharp burst events, such as chiseling, an accurate detection is crucial. For drilling and sawing, the actions span over a longer period of time, as visible in the waveform plot in Figure 3. Therefore, a relaxed criterion for the detection of the onset is tolerable. The detection of drilling events presented a more challenging problem which can mainly be attributed to the fact that the surgeons turned on the device in the beginning of the sequence and kept drilling into the plastic bone while the power tool was running, resulting in only very subtle sound differences between drilling in the air and drilling into the rather soft plastic bone. To mitigate this issue, beam steering or noise canceling could improve the results, which is, however, beyond the scope of the presented work.
The event localization module proved that the fusion of audio-visual information is an effective approach to detect regions of interest with minimal computational overhead. While this method can directly be employed to inform down-stream tasks, such as scene graph generation or instrument tracking, we also see potential as an initial filtering stage that allows more computationally intensive methods to focus solely on the regions of interest. Despite the simplicity of our approach, $84\%$ of the events were successfully localized with a 3D bounding box intersection over union of at least $0.1$. This accuracy is enough to generate semantic scene states, i.e., scene graphs for surgical context understanding. Still, the proposed learning-free baseline and the low resolution of the acoustic camera limit the localization accuracy, as indicated by the steep decrease in recall under stricter intersection over union thresholds. For tasks with higher localization requirements, our localization baseline could be replaced with a deep learning-based approach to enable the extraction of more descriptive multi-modal features, which may enable a more precise sound source localization. An additional potential direction for future research is the projection of more informative acoustic features by evaluating advanced beamforming algorithms. In this context, our current implementation uses a standard delay-and-sum beamforming algorithm that is part of the commercial software shipped with the acoustic camera used within this study.
Our implementation has a latency of about 250ms250\>ms for the event detection and about 100ms100\>ms for the localization stage (measured on a downsampled point cloud with $\frac{1}{64}$ of the original size). Additional optimizations such as GPU-based acceleration and precompiled methods could further decrease these results. While we cannot measure the latency of the beam-forming algorithm due the absence of an API in the closed-source software, comparable methods take around 200ms200\>ms [jansen2025delay]. Our current implementation only supports offline processing for the same reason. Future works should address these limitations.
The exploratory character of our experiments inherently comes with some limitations. The presented dataset is limited in scale and diversity, containing sequences of three representative surgical action, and does not represent the full complexity of a real surgery. Still, our preliminary results show the potential of our novel multimodal audio-visual surgical scene representation and motivate further research to investigate various aspects for improvement, e.g., a wider range of surgical activity, advanced beamforming algorithms, high-quality 3D scene representations and their enrichment via the integration of additional low-level perception algorithms. During the experiment, we used a plastic bone and foam soft tissue model that was moved frequently to introduce variation. The robustness of the proposed learning-free localization approach should be subject to future studies that should evaluate realistic patient models, as well as obtaining larger amounts of data with more variation.
We employ a single-view point cloud as the dynamic 3D visual representation of the surgical scene, which is the basis for mapping the localized acoustic activity. While the ZED2i RGB-D camera produces high-quality point clouds, the preservation of fine details and completeness of the scene is limited. To obtain a higher-quality visual representation, future research should utilize multi-view reconstructions to reduce occlusions and investigate alternative scene representations such as dynamic gaussian splats.
Acoustic information has been shown to have great potential for the analysis of surgical activity in related work, however, acoustic signal analysis is inherently limited to actions that generate audible sounds. While this makes the approach highly effective in sound-rich fields like orthopedics, its application to other surgical sub-fields may be more challenging. However, the proof-of-concept presented in this work does not fully utilize the high sensitivity of the acoustic camera. Especially for hard-to-detect sounds, acoustic beam steering can be a highly promising solution to improve the sound quality for acoustics-based downstream tasks.
6 Conclusion
In summary, this work introduces a novel framework for surgical scene understanding that fuses visual and acoustic information into a unified 4D audio-visual digital representation. By combining RGB-D–based dynamic point clouds with sound localization information obtained from a phased microphone array, and by incorporating transformer-based acoustic event detection, our approach not only enhances contextual awareness in surgical environments but also lays the groundwork for future intelligent and autonomous systems capable of richer, data-driven surgical scene interpretation.
Declarations
Acknowledgements This work has been funded by the Swiss National Foundation under the funding scheme ”Spark” with grant number 228813. The experiments were supported by the OR-X - a Swiss national research infrastructure for translational surgery and associated funding by the University Hospital Balgrist. Imaging was performed with support of the Swiss Center for Musculoskeletal Imaging (SCMI). We thank Dr. Matthias Brechbühl from Norsonic Brechbühl AG and Carsten Hessenius, Andy Meyer, and Michael Markus Ackermann from GFaI e.V. and gfai tech GmbH for their generous support.
Competing Interests The authors have no relevant financial or non-financial interests to disclose.
Ethics Approval Since this study did not involve human or animal subjects, ethics approval is not applicable.
Informed Consent As this research did not involve human participants, informed consent is not applicable.